# The exact travelling wave solutions of a KPP equation

### v2

@@ III The ODE which doesn’t contain explicitly the independent variable
Equation (5) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea kogan3 ; kogan4 ; kogan5 to consider $u$ as the new independent variable and
@@
Now let us be more specific and consider $F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right)$, where $n>0$ and $a\geq 0$ are some constants, so Eqs. (5) and (9) respectively take the form
@@
For the Fisher equation we obtain the well known result ablo
@@
In Appendix A we’ll see how the results (12) - (15) can be generalized in the case when advection is taken into account in addition to diffusion.
@@
and the general integral is $Y=CX+X\ln X$. In order to construct these curves, we can express $X$ and $Y$ in terms of an auxiliary variable by putting $X=e^{\theta}$, which gives $Y=Ce^{\theta}+\theta e^{\theta}$. When $\theta$ approaches $\infty$- ∞, $X$ and $Y$, and consequently $x$ and $y$, approach zero, and the origin is again a focus. Note that if for physical reasons $u$ should be strictly nonnegative, the focus solutions (and hence the value of speed $c$ for which $c^{2}\leq 4F(0)$) are unacceptable.
@@ Appendix A Generalized Burgers–Huxley equation
where $\lambda$ is a constant. (The Burgers–Huxley equation is known in various fields of applied mathematics. For example, it describes transport processes in systems when diffusion and convection are equally important and nonlinear reaction–diffusion processes too.)
@@
Appendix B Perturbation series

### v3

@@ III The ODE which doesn’t contain explicitly the independent variable
Equation (5) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea benguria ; kogan3 ; kogan4 ; kogan5 to consider $u$ as the new independent variable and
@@
Now let us be more specific and consider
$F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (11)
where $n>0$ and $a\geq 0$ are some constants, so Eqs. (5) and (9) respectively take the form
@@
(An alternative way to obtain (16) was presented in our previous publication kogan4 .)
For the Fisher equation from (16) follows the well known result ablo
@@
In Appendix A we’ll see how the results (13) - (16) can be generalized in the case when advection is taken into account in addition to diffusion. In Appendix C we’ll analyze the solution of (12b) (for $n=1$) for $c$ close to $c_{0}$.
@@
and the general integral is $Y=CX+X\ln X$. In order to construct these curves, we can express $X$ and $Y$ in terms of an auxiliary variable by putting $X=e^{\theta}$, which gives $Y=Ce^{\theta}+\theta e^{\theta}$. When $\theta$ approaches $\infty$- ∞, $X$ and $Y$, and consequently $x$ and $y$, approach zero, and the origin is again a focus.
Note that if for physical reasons $u$ should be strictly nonnegative, the focus solutions (and hence the value of speed $c$ for which $c^{2}\leq 4F(0)$) are unacceptable. (Apart from recovering this well known property, we in this paper will say nothing about the front propagation speed selection, an issue which is both important and complicated (saarloos ; ma .)
@@
Acknowledgements.
We are grateful to G. G. Contreras, A. Gupta, M. Ma, B. Malomed and C. Ou for the insightful discussions.
@@ Appendix A Generalized Burgers–Huxley equation
where $\lambda$ is a constant. The Burgers–Huxley equation is known in various fields of applied mathematics. For example, it describes transport processes in systems when diffusion and convection are equally important and nonlinear reaction–diffusion processes too. (An interesting further generalization of the model see in gupta .)
@@
Appendix B Perturbation theory I
@@
Appendix C Perturbation theory II
In Section III we have solved Eq. (12b) for $c=c_{0}$. Now let us study the solution for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series. To make life simpler $n=1$, we’ll consider only the case $n=1$. Presenting the solution as
$p(u)=p_{0}(u)+\frac{\delta c}{\sqrt{2}}\delta p(u),$ (49)
for $\delta p(u)$ in the lowest with respect to $\delta c$ order we obtain equation
$u(u-1)\delta p_{u}+\left(2u-1+c_{0}\right)\delta p+u(u-1)=0.$ (50)
Integrating (50) we obtain
[…]
where
[…]
and $C$ is an arbitrary constant. The lower limits of the integrals in (51) and (52) can be taken arbitrary. Calculating the integral in (52) we get
$\delta p(u)=\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\left(C-\int\frac{\left(1-u\right)^{c_{0}+1}du}{u^{c_{0}-1}}\right).$ (53)
It will be convenient to choose the lower limit in the in integral in (53) to be equal to zero. After that, for $\delta p$ to remain finite when $u\to 0,1$ we should choose $C=0$, thus obtaining
$\delta p(u)=-\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\int_{0}^{u}\frac{\left(1-u\right)^{c_{0}+1}du}{u^{c_{0}-1}}.$ (54)
A pleasant surprise is that $\delta p(u)$, as given by (54), goes to zero when $u\to 0,1$. (We couldn’t be sure in that just looking at Eq. (50), because the coefficient before $\delta p_{u}$ in the equation goes to zero in the above mentioned limits, hence one can doubt the validity of the linear approximation with respect to $\delta c$ in the vicinity of $u=0$ and $u=1$. )